\begin{thebibliography}{10}

\bibitem{BerndtR}
B.~C. Berndt.
\newblock {\em Ramanujan's Notebooks: Part IV}.
\newblock Springer--Verlag, 1993.

\bibitem{DudekCubes}
A.~W. Dudek.
\newblock An explicit result for primes between cubes.
\newblock arXiv:1401.4233v1, January 2014.

\bibitem{DudekPlatt}
A.~W. Dudek and D.~J. Platt.
\newblock Solving a curious inequality of {R}amanujan.
\newblock Preprint available at arXiv:1407.1901.

\bibitem{DusThesis}
P.~Dusart.
\newblock Autour de la fonction qui compte le nombre de nombres premiers, 1998.
\newblock PhD thesis, Universit\'{e} de Limoges.

\bibitem{Dusart}
P.~Dusart.
\newblock Estimates of some functions over primes without {R}.{H}.
\newblock arXiv:1002.0442v1, 2010.

\bibitem{Faber}
L.~Faber and H.~Kadiri.
\newblock New bounds for $\psi(x)$.
\newblock {\em To appear in Math. Comp.}, October 2013.
\newblock Preprint available at arXiv: 1310.6374v1.

\bibitem{Ford}
K.~Ford.
\newblock Vinogradov's integral and bounds for the {R}iemann zeta function.
\newblock {\em Proc. London Math. Soc.}, 85(3):565--633, 2002.

\bibitem{FordInt}
K.~Ford.
\newblock Maximal collections of intersecting arithmetic progressions.
\newblock {\em Combinatorica}, 23(2):263--281, 2003.

\bibitem{Ford04}
K.~Ford.
\newblock A strong form of a problem of {R}. {L}. {G}raham.
\newblock {\em Canad. Math. Bull.}, 47(3):358--368, 2004.

\bibitem{HST}
L.~Hajdu, N.~Saradha, and R.~Tijdeman.
\newblock On a conjecture of {P}omerance.
\newblock {\em Acta Arith.}, 155(2):175--184, 2012.

\bibitem{Helfgott}
H.~Helfgott.
\newblock Major arcs for {G}oldbach's problem.
\newblock arXiv:1305.2897v2, 2013.

\bibitem{Kadiri}
H.~Kadiri.
\newblock Une r\'{e}gion explicite sans z\'{e}ros pour la fonction $\zeta$ de
  {R}iemann.
\newblock {\em Acta Arith.}, 117(4):303--339, 2005.

\bibitem{Kotnik}
T.~Kotnik.
\newblock The prime-counting function and its analytic approximations.
\newblock {\em Adv. Comput. Math.}, 29(1):55--70, 2008.

\bibitem{NymNic}
T.~R. Nicely and B.~Nyman.
\newblock New prime gaps between $10^{15}$ and $5\times 10^{16}$.
\newblock {\em J. Integer Seq.}, 6(3):1--6, 2003.

\bibitem{Plattarxiv}
D.~J. Platt.
\newblock Computing $\pi(x)$ analytically.
\newblock arXiv:1203.5712v3, October 2013.
\newblock To appear in \textit{Math. Comp.}

\bibitem{PomC}
C.~Pomerance.
\newblock A note on the least prime in an arithmetic progression.
\newblock {\em J. Number Theory}, 12:218--223, 1980.

\bibitem{RamRum}
O.~Ramar\'{e} and R.~Rumely.
\newblock Primes in arithmetic progressions.
\newblock {\em Math. Comp.}, 65(213):397--425, 1996.

\bibitem{RamSao}
O.~Ramar\'{e} and Y.~Saouter.
\newblock Short effective intervals containing primes.
\newblock {\em J. Number Theory}, 98:10--33, 2003.

\bibitem{RS1}
J.~B. Rosser and L.~Schoenfeld.
\newblock Approximate formulas for some functions of prime numbers.
\newblock {\em Illinois J. Math.}, 6:64--94, 1962.

\bibitem{RS2}
J.~B. Rosser and L.~Schoenfeld.
\newblock Sharper bounds for the {C}hebyshev functions $\theta(x)$ and
  $\psi(x)$.
\newblock {\em Math. Comp.}, 29(129):243--269, 1975.

\bibitem{Schoenfeld}
L.~Schoenfeld.
\newblock Sharper bounds for the {C}hebyshev functions $\theta(x)$ and
  $\psi(x)$, {II}.
\newblock {\em Math. Comp.}, 30(134):337--360, 1976.

\bibitem{Togbe}
A.~Togb\'{e} and S.~Yang.
\newblock Proof of the ${P}$-integer conjecture of {P}omerance.
\newblock {\em J. Number Theory}, 140:226--234, 2014.

\end{thebibliography}

\documentclass[12pt]{article}
\title{Updating the error term in the prime number theorem}
\author{Tim Trudgian\footnote{Supported by Australian Research Council DECRA Grant DE120100173.}\\ Mathematical Sciences Institute\\ The Australian National University, ACT 0200, Australia\\ timothy.trudgian@anu.edu.au }

\usepackage{url}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{booktabs}
\newtheorem{thm}{Theorem}
\newtheorem{ex}{Example}
\newtheorem{cor}{Corollary}
\newtheorem{lem}{Lemma}

\begin{document}

\maketitle

\begin{abstract}
\noindent An improved estimate is given for $|\theta(x) -x|$, where $\theta(x) = \sum_{p\leq x} \log p$. Four applications are given: the first to arithmetic progressions that have points in common, the second to primes in short intervals,  the third to a conjecture by Pomerance, and the fourth to an inequality studied by Ramanujan.
\\
\\
Key words: Prime number theorem; Chebyshev functions; Ramanujan's inequality\\ \\
AMS Codes: 11M06, 11N05\\ \\
Dedicated to MG Johnson, RJ Harris, PM Siddle, and NM Lyon, all of whom enabled me to work two extra days on this article.
\end{abstract}

\section{Introduction}
One version of the prime number theorem is that $\theta(x) \sim x$, where $\theta(x) = \sum_{p\leq x} \log p$. Several applications call for an explicit estimate on the error $\theta(x) - x$. Schoenfeld, \cite[Thm 11]{Schoenfeld} proved that for
\begin{equation}\label{ep}
\epsilon_{0}(x) = \sqrt{\frac{8}{17 \pi}} X^{1/2}e^{-X}, \quad X = \sqrt{(\log x)/R_{0}}, \quad R_{0}= 9.6459, 
\end{equation}
the following inequality holds
$$|\theta(x) - x| \leq x \epsilon_{0}(x), \quad (x \geq 101).$$ 
The pair of numbers $(R_{0}, 17)$ in (\ref{ep}) is particularly interesting. These arise from \cite[Thm 1]{RS2}, namely, the theorem that 
\begin{equation}\label{zero}
\textrm{$\zeta(s)$ has no zeroes in the region} \; \sigma \geq 1- \frac{1}{R \log |\frac{t}{B}|}, \quad (t\geq t_{0})
\end{equation}
for $R= R_{0}$, $B= 17$ and $t_{0} = 21$. Ramar\'{e} and Rumely \cite[p.\ 409]{RamRum} proved (\ref{zero}) with $(R, B, t_{0}) = (R_{0}, 38.31, 1000)$; Kadiri \cite{Kadiri} proved (\ref{zero}) with $(R, B, t_{0}) = (5.69693, 1, 2)$.

A meticulous overhaul of Schoenfeld's paper would be required to furnish a `general' version of (\ref{zero}), that is, one in which $B$ and $R$ are chosen for maximal effect. This article does not attempt such an overhaul. Rather, forcing $B$ to be $17$ in (\ref{zero}) means that many of the numerical estimations in Schoenfeld's article can be let through to the keeper. With $B= 17$ one can obtain admissible values of $R$ and $t_{0}$ in (\ref{zero}) as follows.

Let the Riemann hypothesis be true up to height $H$: by Platt \cite{Plattarxiv} we have $H= 3.061\times 10^{10}$. Let $\rho$ represent a non-trivial zero of $\zeta(s)$ with $\rho = \beta + i\gamma$. Using Kadiri's result we see that
\begin{equation*}
\beta \leq 1 - \frac{1}{5.69693 \log t} \leq 1 - \frac{1}{R \log |\frac{t}{17}|}
\end{equation*}
provided that 
$$t\geq \exp\left\{\frac{R \log 17}{R - 5.69693}\right\}.$$
Set $H = \exp\{R \log 17/(R - 5.69693)\}$, whence we may take $R = 6.455$. We conclude that there are no zeroes in 
\begin{equation}\label{zero2}
\sigma \geq 1 - \frac{1}{6.455 \log|\frac{t}{17}|}, \quad (t\geq 24).
\end{equation}
This enables us to prove good bounds for $\theta(x) -x$ and for $\psi(x) -x$, where $\psi(x) = \sum_{p^{m}\leq x} \log p$, as indicated in the following theorem.
\begin{thm}\label{TSwan}
Let 
\begin{equation*}\label{ep2}
\epsilon_{0}(x) = \sqrt{\frac{8}{17 \pi}} X^{1/2}e^{-X}, \quad X = \sqrt{(\log x)/R}, \quad R = 6.455.
\end{equation*}
Then
\begin{equation*}
\begin{split}
|\theta(x) - x| &\leq x \epsilon_{0}(x), \quad (x \geq 149)\\
|\psi(x) - x| &\leq x \epsilon_{0}(x), \quad (x \geq 23).
\end{split}
\end{equation*}
\end{thm}
Throughout Schoenfeld's paper numerous bounds on $x$ are imposed, where $X = \sqrt{(\log x)/R_{0}}$. Fortunately, for our purposes, all of these arise from bounds imposed on $X$. For example, the first bound in \cite[(7.30)]{Schoenfeld} requires $X \geq 17/2\pi$. With our value of $R$ we need $\log x \geq 48$ compared with Schoenfeld's requirement  $\log x \geq 71$.
Making these slight changes throughout pp.\ 342-348 of \cite{Schoenfeld} we find that
\begin{equation}\label{in}
|\psi(x) - x|,  |\theta(x) - x| \leq x \epsilon_{0}(x), \quad (\log x \geq 1163).
\end{equation}
In order to prove Theorem \ref{TSwan} we cover small values of $x$ following the approach on pp.\ 348-349 of \cite{Schoenfeld} but using the superior bounds on $|\psi(x) -x|$ as given by Faber and Kadiri \cite{Faber}.
We make use of equation ($5.3^{*}$) in \cite{Schoenfeld}, namely, 
\begin{equation*}
\psi(x) - \theta(x) < 1.001093 x^{1/2} + 3x^{1/3}
 \leq A(x_{0}) x, \quad (x\geq x_{0})
\end{equation*}
where $A(x_{0}) = 1.001093 x_{0}^{-1/2} + 3x_{0}^{-2/3}.$
For $e^{25} \leq x \leq e^{45}$ we have, by \cite[Table~3]{Faber},
\begin{equation*}
|\psi(x) - x|, \quad |\theta(x) - x| \leq (A(e^{25}) +4.9\times 10^{-5}) \frac{x\epsilon_{0}(x)}{\epsilon_{0}(e^{45})} \leq 0.003 x\epsilon_{0}(x).
\end{equation*}
Now for $e^{45} \leq x \leq e^{1163}$ we have
\begin{equation*}
|\psi(x) - x|, \quad |\theta(x) - x| \leq (A(e^{45}) +1.1\times 10^{-8}) \frac{x\epsilon_{0}(x)}{\epsilon_{0}(e^{1162})} \leq 0.006 x\epsilon_{0}(x).
\end{equation*}
Hence (\ref{in}) is true for all $x\geq e^{25}$. For $x< e^{25}$ note that $\epsilon_{0}(x)$ increases for $X<\frac{1}{2}$ and decreases thereafter. Therefore
\begin{equation}\label{last}
\epsilon_{0}(x) \geq \min\{\epsilon_{0}(2), \epsilon_{0}(e^{25})\} \geq 0.075.
\end{equation}
Theorem 10 in \cite{RS1} gives $\theta(x) > 0.93x$ for $x\geq 599$. This, combined with (\ref{last}), shows that 
\begin{equation}\label{had}
\theta(x) -x > -x\epsilon_{0}(x), \quad (x\geq 599).
\end{equation}
Since $\psi(x) \geq \theta(x)$ the inequality in (\ref{had}) also holds with $\psi(x)$ in place of $\theta(x)$.
Using $\psi(x) \leq 1.04 x$ (see Theorem 12 in \cite{RS1}) and (\ref{last}) gives
\begin{equation*}\label{mu}
\theta(x) \leq \psi(x) \leq 1.04 x < x+ x\epsilon_{0}(x), \quad (2\leq x \leq e^{25}).
\end{equation*}
All that remains is to verify (\ref{had}) and the analogous inequality for $\psi(x)$ for values of $x\leq 599$ --- a computational dolly. 

\section{The difference $\pi(x) - \textrm{li}(x)$}
Let $\pi(x)$ denote the number of primes not exceeding $x$ and $\textrm{li}(x)$ denote the logarithmic integral, namely
$$\textrm{li}(x) = \lim_{\epsilon\rightarrow 0^{+}} \left(\int_{0}^{1- \epsilon} \frac{dt}{\log t} + \int_{1+ \epsilon}^{x} \frac{dt}{\log t}\right).$$
Concerning the difference $\pi(x) - \textrm{li}(x)$ we have
\begin{equation}\label{pill2}
|\pi(x) - \textrm{li}(x)| \leq 0.4394 \frac{x}{(\log x)^{3/4}} \exp(-\sqrt{(\log x)/9.696}), \quad (x\geq 59),
\end{equation}
due to Dusart \cite[Thm 1.12]{DusThesis}\footnote{There is also the result of Ford \cite{Ford} 
$$\pi(x) - \textrm{li}(x) = O( x \exp\{-0.2098 (\log x)^{3/5} (\log\log x)^{-1/5}\}).$$ It appears that this result has not been made explicit.}. 
Good bounds on $\pi(x) - \textrm{li}(x)$ can be obtained from good bounds on $\theta(x) -x$, since
\begin{equation}\label{pi}
\begin{split}
\pi(x) - \textrm{li}(x) &= \frac{\theta(x)}{\log x} + \int_{2}^{x}\frac{\theta(t)}{t \log^{2} t}\, dt - \int_{2}^{x} \frac{dt}{\log t} - \textrm{li}(2)\\
&=  \frac{\theta(x)-x}{\log x} +  \frac{2}{\log 2} +  \int_{2}^{x}\frac{\theta(t)-t}{t \log^{2} t}\, dt - \textrm{li}(2).
\end{split}
\end{equation}
Using Theorem \ref{TSwan} we can prove
\begin{thm}\label{TLi}
\begin{equation*}
|\pi(x) - \textrm{li}(x)| \leq 0.2795 \frac{x}{(\log x)^{3/4}} \exp\left(-\sqrt{\frac{\log x}{6.455}}\right), \quad (x\geq 229).
\end{equation*}
\end{thm}
\begin{proof}
We split up the range of integration in (\ref{pi}) so that $\int_{2}^{x} = \int_{2}^{x_{0}} + \int_{x_{0}}^{x} = I_{1} + I_{2}$ for some $x_{0}\geq 149.$ To estimate $I_{2}$ we use Theorem \ref{TSwan} and consider
$$g(t) = \frac{t \epsilon_{0}(t)}{(\log t)^{\alpha + \frac{1}{4}}}.$$
The value of $\alpha$ in the expression for $g(t)$ must be less than $7/4$. Following Dusart we choose $\alpha = 7/5$, whence it is easy to verify that $\epsilon_{0}(t)/(\log t)^{2} < g'(t)$ for all $t\geq 149$.

To estimate $I_{1}$ we invoke \cite[Thm 19]{RS1}
$$ \theta(t) < t, \quad t<10^{8}.$$
Interchanging summation and integration we have
\begin{equation*}\label{dust}
\int_{2}^{x_{0}} \frac{\theta(t)}{t\log^{2} t}\, dt = \int_{2}^{x_{0}} \frac{\sum_{p\leq t}\log p}{t \log^{2} t}\, dt =  \sum_{p\leq x_{0}} \log p \int_{p}^{x_{0}} \frac{dt}{t \log^{2} t} = \pi(x_{0}) - \frac{\theta(x_{0})}{\log x_{0}}.
\end{equation*}
Therefore (\ref{pi}) becomes
\begin{equation*}\label{kite}
\begin{split}
|\pi(x) - \textrm{li}(x)| \leq &\frac{x \epsilon_{0}(x)}{\log x} + \frac{x\epsilon_{0}(x)}{(\log x)^{33/20}} + \frac{2}{\log 2} - \textrm{li}(2)  -\frac{x_{0}\epsilon_{0}(x_{0})}{(\log x_{0})^{33/20}} \\
&+ \int_{2}^{x_{0}} \frac{dt}{\log^{2} t} -\pi(x_{0}) + \frac{\theta(x_{0})}{\log x_{0}}.
\end{split}
\end{equation*}
We may choose $x_{0}$ in (\ref{pi}) subject to $149\leq x_{0}\leq 10^{8}$. Choosing $x_{0} = 10^{8}$ shows, in less than 3 minutes  using \textit{Mathematica} on a 1.8GHz laptop, that
\begin{equation}\label{pi2}
|\pi(x) - \textrm{li}(x)| \leq \frac{x \epsilon_{0}(x)}{\log x} + \frac{x\epsilon_{0}(x)}{(\log x)^{33/20}}  \leq 1.151 \frac{x \epsilon_{0}(x)}{\log x},
\end{equation}
for $x\geq 10^{8}$. For smaller $x$ we note that, by Kotnik \cite{Kotnik} $\pi(x) < \textrm{li}(x)$ for $2\leq x \leq 10^{14}$. Therefore
\begin{equation}\label{picheck}
\max_{x\in[p_{k}, p_{k+1})} |\pi(x) - \textrm{li}(x)| = \max_{x\in[p_{k}, p_{k+1})} \textrm{li}(x) - \pi(x) \leq \textrm{li}(p_{k+1}) - k.
\end{equation}
Using (\ref{picheck}) we verify (\ref{pi2}) for all $x\geq p_{k}$ with $k \geq 48$, which is equivalent to $x\geq 229$, which proves the theorem.
\end{proof}

\section{Applications}
We now present four applications of Theorem \ref{TSwan} and \ref{TLi}. We stress that explicit results of this nature have many uses throughout the literature; our list of four applications is by no means exhaustive. One striking example of this applicability is Helfgott's proof of the ternary Goldbach conjecture \cite{Helfgott}. In \cite[\S 7]{Helfgott} Helfgott makes frequent use of estimations for the number of primes in short intervals and the size of the Chebyshev functions. 
\subsection{Intersecting arithmetic progressions} 
Let $N_{t}(k)$ denote the maximum number of distinct arithmetic progressions of $k$ numbers such that any pair of progressions has $t$ members in common. Ford \cite{FordInt} considers the following example.

\begin{ex}
For $1\leq i < j \leq k$, let $B_{ij}$ be the arithmetic progression the $i$th element of which is 0, and the $j$th element of which is $k!$.
\end{ex}

Ford shows, in Theorem 3 of \cite{FordInt}, that for all $k\geq 10^{8000}$, $N_{2}(k) = k(k-1)/2$, and that every configuration of $k(k-1)/2$ arithmetic progressions with 2 points in common is equivalent (up to translations and dilations) to the arithmetic progression in Example 1. We are able to use Theorem \ref{TSwan} to prove
\begin{cor}\label{ford}
For $k \geq 10^{4848}$ we have $N_{2}(k) = k(k-1)/2$ and that every configuration of $k(k-1)/2$ arithmetic progressions with 2 points in common is equivalent to the arithmetic progression in Example 1.
\end{cor}
\begin{proof}
Analogous to Lemmas 3.3 and 3.4 in \cite{FordInt} we can show that for $k\geq e^{280}$ there is always a prime in the interval $[k, k+a]$ where \begin{equation}\label{adef}
a = 0.56 k(\log k)^{1/4} \exp(-\sqrt{(\log k)/6.455}).
\end{equation} 
We now follow the proof of Theorem 3 in \cite{FordInt} using our (\ref{adef}) in place of his $a =  0.44 k(\log k)^{1/4} \exp(-0.321979\sqrt{\log k}).$
\end{proof}
It is worthwhile to remark that Corollary \ref{ford} could be improved if the method in \cite{Ford04} were made explicit. However, it seems unlikely that one could reduce the bound on $k$ in Corollary \ref{ford} to a height below which direct computation could be carried out.

\subsection{Primes in short intervals}
Various results have been proved about the existence of a prime in a short interval $[x, x+ f(x)]$ where $f(x) = o(x)$. For example, Dusart \cite[Prop.\ 6.8]{Dusart} has shown that there exists a prime in the interval $[x, x + \frac{x}{25\log^{2}x}]$ whenever $x\geq 396738$. We improve this in
\begin{cor}\label{primes}
For all $x\geq 2898239$ there is a prime in the interval
$$ \left[x, x\left(1 + \frac{1}{111\log^{2}x}\right)\right].$$
\end{cor}
We first prove the following
\begin{lem}\label{lemonly}
For $x\geq e^{35}$ we have
\begin{equation}\label{thmain}
|\theta(x)- x| \leq \frac{0.0045x}{\log^{2} x}.
\end{equation}
\end{lem}
\begin{proof}
Using ($5.3^{*}$) of \cite{Schoenfeld} we have
\begin{equation}\label{ee}
\frac{|\theta(x) -x|\log^{2} x}{x} \leq \left(\frac{|\psi(x) -x|}{x} \log^{2}x + \frac{1.001093 \log^{2} x}{\sqrt{x}} + \frac{3\log^{2}x}{x^{2/3}}\right)= B(x),
\end{equation}
say.
According to Table 3 in \cite{Faber}, $|\psi(x) - x|\leq 7.4457\times 10^{-7}$ for $x\geq e^{35}$, whence $B(x)$ is bounded above by
$$ \left(7.4457\times 10^{-7} \log^{2}x + \frac{1.001093 \log^{2} x}{\sqrt{x}} + \frac{3\log^{2}x}{x^{2/3}}\right)\bigg|_{x = e^{75}}, \quad x\in[e^{35}, e^{75}],$$
which is bounded above by 0.0042. We continue in this way, using intervals of the form $[e^{a}, e^{b}]$, Faber and Kadiri's bounds at $e^{a}$ and evaluating $B(x)$ at $x=e^{b}$. The results are summarised below in Table \ref{table}. 

\begin{table}[ht]
\caption{Bounding $\theta(x) - x$}
\label{table}
\centering
\begin{tabular}{lc}
\hline\hline
Interval & Bound on $B(x)$ in (\ref{ee})\\[0.5ex]\hline
$[e^{35}, e^{75}]$ & 0.0042\\
$[e^{75}, e^{1500}]$ & 0.0037\\
$[e^{1500}, e^{2000}]$ & 0.0038\\
$[e^{2000}, e^{2500}]$ & 0.0045\\
$[e^{2500}, e^{3000}]$ & 0.0044\\
$[e^{3000}, e^{4000}]$ & 0.0036\\
    \hline\hline
  \end{tabular}
\end{table}
Taking the maximum entry in the right-hand column of the table proves Lemma \ref{lemonly} for $e^{35}\leq x \leq e^{4000}$. When $x\geq e^{4000}$ we use Theorem \ref{TSwan}. This completes the proof of the lemma.
\end{proof}
Note that one could refine this result by taking more intermediate steps in the argument. For example one could use the interval $[e^{2000}, e^{2100}]$ to try to reduce the bound of $0.0045$. We have not pursued this since the entry $e^{2100}$ is not in Table 3 in \cite{Faber} and, while it could be calculated, the above lemma is sufficient for our purposes.

We now use Lemma \ref{lemonly} to exhibit primes in short intervals. Indeed, for $x\geq e^{35}$ Lemma \ref{lemonly} shows that
$$\theta\left\{x\left(1 + \frac{1}{c\log^{2} x}\right)\right\} - \theta(x)$$
is positive provided that $c\leq 111.1107\ldots$. Taking $c=111$ we conclude that there is always a prime in the interval $[x, x(1+ 1/(111 \log^{2} x))]$ whenever $x\geq e^{35}$.
This establishes Corollary \ref{primes} when $x\geq e^{35} \approx 1.58\times 10^{15}$. Rather than perform the herculean, if not impossible, feat of examining all those $x< e^{35}$ we proceed as follows.

Suppose that $p_{n+1} - p_{n}\leq X_{1}$ for all $p_{n} \leq x_{1}$, where $x_{1} \geq e^{35}$. That is, the maximal prime gap of all primes up to $x_{1}$ is at most $X_{1}$. Therefore $p_{n+1} \leq p_{n} + X_{1}$ which will be less that $p_{n}(1+ \frac{1}{111 \log^{2}p_{n}})$ as long as
\begin{equation}\label{p1}
\frac{p_{n}}{\log^{2}p_{n}} \geq 111 X_{1}.
\end{equation}
If (\ref{p1}) holds for all $y_{1} \leq p_{n} \leq x_{1}$ we can conclude that Corollary \ref{primes} holds for all $x\geq y_{1}$. If $y_{1}$ is still too high for a direct computation over all integers less than $y_{1}$, then we may play the same game again, namely: find an $x_{2}\geq y$ such that $p_{n+1} - p_{n} \leq X_{2}$.

Nyman and Nicely \cite[Table 1]{NymNic} show that one may take $x_{1} = 1.68\times 10^{15}$, which is greater than $e^{35}$, and $X_{1} = 924$. It is easy to verify that (\ref{p1}) holds for all $p_{n} \geq 3.05\times 10^{7}$.
We can now check, relatively swiftly that the maximal prime gap for $p_{n} < 3.06\times 10^{7}$ is $210$. We may now verify Corollary \ref{primes} for all $x\geq 5.63\times 10^{6}$. Two more applications of this method, using the fact that the maximal prime gap for $p_{n} < 5.7\times 10^{6}$ is $159$, and for $p_{n} < 4\times 10^{6}$ is $148$ we see that Corollary \ref{primes} is true for all $x\geq 3.8\times 10^{6}$.

We now examine $x \leq 3.8\times 10^{6}$. An exhaustive search took less than two minutes on \textit{Mathematica} --- this completes the proof of Corollary \ref{primes}.

There are several ways in which this result could be improved. Extending the work done by Nyman and Nicely \cite{NymNic} makes a negligible difference to the choice of $c$. Probably the best plan of attack is reduce the size of the coefficient in Lemma \ref{lemonly}. For example, if the coefficient in (\ref{thmain}) were reduced to $0.0039$ we could take $c=128$.

Finally, the result in Corollary \ref{primes} ought to be compared with the sharpest known result for a different short interval. Ramar\'{e} and Saouter \cite[Table 1]{RamSao} proved that there is always a prime in the interval
$$(x(1-\Delta^{-1}), x],\quad \Delta = 212215384, \quad x\geq e^{150}.$$
Corollary \ref{primes} improves on this whenever $x\geq 3.2\times 10^{600} \approx e^{1383}$. Although this value of $x$ is large by anyone's standards, it appears that Corollary \ref{primes} could be useful in searching for primes between cubes --- see \cite{DudekCubes}.

\subsection{A conjecture by Pomerance}
Consider numbers $k>1$ for which the first $\phi(k)$ primes coprime to $k$ form a reduced residue system modulo $k$. Following the lead of Hajdu, Saradha and Tijdeman \cite{HST}, we call such an integer $k$ a \textit{P-integer}. For example $12$ is a $P$-integer and $10$ is not since
$$\{5,7,11,13\} \equiv \{5, 7, 11, 1\}, \quad \quad \{3,7,11,13\} \equiv\{ 3,7,11,3\},$$
and, whereas the first is a reduced residue system, the second is not.
From \cite[Thm 2]{PomC} Pomerance deduced that there can be only finitely many $P$-integers. Hajdu, Saradha and Tijdeman \textit{[op.\ cit.]} proved, \textit{inter alia}, that if $k$ is a $P$-integer such that $k> 30$ then
$ 10^{11}< k < 10^{3500}.$
As noted by Hajdu, Saradha and Tijdeman, one may improve (\ref{pill2}) by using the zero-free region proved by Kadiri, that is, by using our Theorem \ref{TSwan}. We do this thereby proving
\begin{cor}\label{T1}
If $k$ is a $P$-integer then $k < 10^{1805}$.
\end{cor}
\begin{proof}
We use Theorem \ref{TLi} instead of Lemma 2.1(iii) in \cite{HST} and proceed as in \cite[\S 5]{HST}. Let $k\geq 10^{1805}$ and define
\begin{equation}\label{fee}
\begin{split}
f_{0}(k) &= \frac{k}{\log k/2} + \frac{k}{\log^{2} k/2} + \frac{1.8 k}{\log^{3} k/2} - \frac{k}{\log k} - \frac{k}{\log^{2}k} - \frac{2.51 k}{\log^{3} k} - \log k,\\
f_{n}(k) &= \frac{k}{4(n+1)\log^{2}(nk +k)} - 1.118 \frac{nk +k}{(\log nk)^{3/4}} \exp( - \sqrt{\log(nk)/6.455}),
\end{split}
\end{equation}
where the constant $1.118$ is four times that appearing in Theorem \ref{TLi}.
Lemma 3.1 in \cite{HST} gives the following
\begin{equation}\label{test1}
\textrm{ $k$ is not a $P$-integer if} \quad f_{0}(k) + \sum_{n=1}^{L} f_{n}(k) >0,
\end{equation}
where $L$ satisfies
$$ L \geq \frac{\log k - \log h(k)}{h(k)} -2, \quad  h(k) = 1.7811\log\log k + 2.51/(\log\log k).$$
When $k\geq 10^{1805}$ we have $L\geq 273$. We verify that the condition in (\ref{test1}) is met for $273\leq L \leq 3800$. We now proceed as in \cite[p.\ 181]{HST} with $3800$ and $k= 10^{1805}$ taking the place of $1500$ and $k= 10^{3500}$ respectively.
\end{proof}
The numbers 1.8 and 2.51 appearing in (\ref{fee}) are worth a mention. These are approximations to the number 2 that appears in the expansion
$$\pi(x) \sim \frac{x}{\log x} + \frac{x}{\log^{2} x} + \frac{2x}{\log^{3} x} + \cdots.$$
Replacing these numbers in (\ref{fee}) by 2, a situation on which one could not possibly improve, makes a negligible difference. Indeed, such a substitution could not improve the bound in Corollary 3 to $k< 10^{1803}$.

It is certainly possible that a refined version of Theorem \ref{TSwan} could resolve completely the Pomerance conjecture. Indeed, using a slightly different approach, Togb\'{e} and Yang have announced in \cite{Togbe} a proof of the conjecture.

\subsection{An equality studied by Ramanujan}
Ramanujan \cite[Ch.\ 24]{BerndtR} proved that
\begin{equation}\label{ram}
\pi(x)^{2} < \frac{e x}{\log x} \pi\left( \frac{x}{e}\right),
\end{equation}
holds for all sufficiently large values of $x$. In a paper to appear, Dudek and Platt \cite{DudekPlatt} have used Theorem \ref{TSwan} to show that, on the Riemann hypothesis, (\ref{ram}) is true for all $x> 38,358,837,682.$ It seems difficult to prove this unconditionally: in this case Dudek and Platt are able to show that (\ref{ram}) is true for all $x\geq \exp(9658)$.

\section{Conclusion}
Theorems \ref{TSwan} and \ref{TLi} could be improved in several ways. First, if one knew that the Riemann hypothesis had been verified to a height greater than $3.061\times 10^{10}$, one could reduce the coefficient in the zero-free region in (\ref{zero2}). Second, one could try to improve Kadiri's zero-free region either by reducing the value of $R$ or by improving the size of $B$ in (\ref{zero}). A higher verification of the Riemann hypothesis has a mild influence on this method of proof.

Third, one may feed any improvements in a numerical verification of the Riemann hypothesis and the zero-free region into Faber and Kadiri's argument, thereby improving the estimate on $\psi(x) - x$. Finally, one may try to overhaul completely Schoenfeld's paper in order to provide a bespoke version of Theorem~\ref{TSwan}. 


\bibliographystyle{plain}
\bibliography{themastercanada}

\end{document}